# Filamentary Ejecta Network in Cassiopeia A Reveals Fingerprints of the Supernova Explosion Mechanism

### v1

@@ 1 Introduction
In this paper, we investigate this issue using the 3D MHD model from Orlando et al. (2021), complemented by high-resolution simulations designed to capture ejecta features at scales approaching those observed by JWST. This model traces the evolution of a neutrino-driven SN explosion, replicating key characteristics of Cas A from the immediate post-core-collapse phase to its current state. In such a way, this approach provides a robust framework to link the observed filaments with specific physical processes active throughout the SN and the remnant evolution. Our study focuses on two main objectives: (i) characterizing the morphological and kinematic properties of the network of filament by comparing the simulated structures with the results from observations, (ii) tracing the ejecta evolution from the initial seconds after the explosion to examine how filamentary features observed today are connected to physical processes and phenomena triggered shortly after core collapse.
@@
The paper is organized as follows: Sect. 2 summarizes our MHD simulation setup and methodology; Sect. 3 presents the results, focusing on the 3D structure of the unshocked ejecta and their origins tracing back to the core-collapse; finally, Sect. 4 summarizes our main conclusions and discusses their broader implications for understanding SN explosion mechanisms and remnant evolution. In Appendix A, we investigate the effects of the spatial resolution on the results; in Appendix B, we evaluate the effects of the magnetic field on the structure of unshocked ejecta; in Appendix C, we describe the content of online multi-media material, offering interactive and dynamic representations of the phenomena discussed in the paper.
2 The 3D MHD model
@@
The output from the 3D SN simulation at $\approx 17.85$ hours after core-collapse served as the initial condition for 3D simulations that track the transition from the SN phase to the SNR phase and the subsequent expansion of the remnant through the CSM (Orlando et al. 2021, 2022). The CSM environment is modeled as a stellar wind with a density profile proportional to $r^{-2}$, normalized to $0.8$ cm-3 at a radius of 2.5 pc (in agreement with Lee et al. 2014), and includes an asymmetric, dense circumstellar shell with its densest region positioned on the near side to the northwest (Orlando et al. 2022). This shell is included to account for significant asymmetries of the reverse shock, which cannot be explained by models assuming a spherically symmetric stellar wind from the progenitor (see Orlando et al. 2022). Recent JWST observations support the existence of this shell, having revealed evidence of a shocked, dense CSM structure around Cas A (the so-called “Green Monster”), consistent with past interactions between the remnant and a dense region of CSM material (Milisavljevic et al. 2024; De Looze et al. 2024; Orlando et al., in preparation). In the MHD simulations, the ambient magnetic field is represented by a “Parker spiral” configuration, resulting from the progenitor star’s rotation and the corresponding outward flow of the stellar wind (Parker 1958; see Orlando et al. 2019a for further details).
The SNR simulations include: (i) energy deposition from radioactive decay, specifically the 56Ni $\rightarrow$ 56Co $\rightarrow$ 56Fe chain, modeled by an internal energy term that assumes local deposition, excluding neutrinos that escape freely, and assuming no $\gamma$-ray leakage from the inner part of the remnant (Orlando et al. 2021); (ii) deviations from ionization equilibrium, estimated through the maximum ionization age in each cell of the spatial domain (Orlando et al. 2015); and (iii) deviations from electron-proton temperature equilibration, with initial electron heating at shock fronts up to $kT=0.3$ keV (Ghavamian et al. 2007), and post-shock ion/electron temperatures determined via Coulomb collisions (Orlando et al. 2015). In this paper, we omitted the impact of cosmic ray back-reaction at shock fronts explored in previous studies (Orlando et al. 2016, 2022). This process primarily influences the dynamics of the forward shock and the structure of the mixing region between the forward and reverse shocks. However, it does not significantly affect the evolution of the unshocked ejecta, which is the primary focus of this work.
@@ 3 Results
Table 2: Masses of ejecta elements (total, shocked, and unshocked) in the W15-IIb-sh-HD+dec-hr model at 351 years. The table also includes the ratio of shocked to unshocked masses. The last three columns provide, for comparison, the shocked mass values inferred from Chandra data analysis of Cas A (Hwang & Laming 2012) and the unshocked mass values inferred from Spitzer data analysis (Laming & Temim 2020).
@@
$(a)$ Data from Hwang & Laming (2012); each element is presented with two values, corresponding to assumed filling factors based on density spike thickness of $2.5^{\prime\prime}$ and $5^{\prime\prime}$; $(b)$ Data from Laming & Temim (2020); $(c)$ The ejecta mass of 44Ti is overestimated by a factor $10-20$ due to a reduced network in the original SN model, which included only 9 of 13 $\alpha$ nuclei up to 56Ni (see Wongwathanarat et al. 2017); $(d)$ The values reported consider the total Fe masses from incomplete Si burning (FeSi) plus from complete Si burning and/or $\alpha$-rich freezout (Feα) (Hwang & Laming 2012); $(e)$ The tracer nucleus 56X represents Fe-group species synthesized in neutron-rich environments as, for instance, in neutrino-heated ejecta (see Wongwathanarat et al. 2017). The table lists the mass of 56Fe, calculated as $M_{\rm Fe}+0.5M_{\rm X}$, and the mass of 56X, estimated as $0.5M_{\rm X}$.
@@ 3.1 The structure of unshocked ejecta at the age of Cas A
From our analysis of the distribution of unshocked species, we found no significant differences between models with or without magnetic fields (see Table 1). Therefore, in the following, we focused on the results of the run W15-IIb-sh-HD+dec-hr, noting any differences observed in the magnetized case (W15-IIb-sh-MHD+dec-hr) where relevant. In Appendix B, we report a selection of results from run W15-IIb-sh-MHD+dec-hr. The effects of magnetic fields, particularly in shaping the structure of the Green Monster, are discussed in a separate study (Orlando et al., submitted to A&A). Unless otherwise stated, the time in this study is measured from the core bounce (see Wongwathanarat et al. 2017), also referred to as the time after core-collapse.
@@
Figure 6: Maps of the filling factor of O-rich ejecta, calculated for cells where the O density exceeds thresholds of 0.1% (upper left panel), 1% (upper right), or 10% (lower left) of the maximum O-rich ejecta density in the unshocked volume. The maps assume the model orientation matching Cas A. The lower-right panel shows the corresponding normalized LoS velocity distributions, derived from the same cells used to derive the filling factor maps.
@@
To investigate further this issue and to facilitate a more quantitative comparison between model predictions and JWST observations, we calculated the filling factor of O-rich ejecta and the corresponding LoS velocity distributions. Figure 6 displays maps of the filling factor, assuming the model is oriented to match Cas A. These maps represent the filling factors computed along the LoS for cells where the O density exceeds specific thresholds ($f_{\rm ths}=$ 0.1%, 1%, or 10%) of the maximum O-rich ejecta density within the entire unshocked volume. Specifically, for each LoS, the filling factor is calculated as the ratio of the volume of cells containing O-rich ejecta above a given density threshold ($V_{\rm O}$) to the total volume of all cells containing unshocked ejecta along the same LoS ($V_{\rm tot}$). This ratio provides a normalized measure of the extent to which the unshocked ejecta volume is occupied by O-rich material, offering valuable insights into its spatial distribution and density contrasts. The lower-right panel of the figure shows the corresponding normalized LoS velocity distributions. These distributions are derived from the same cells used to compute the filling factors and highlight the kinematic properties of the O-rich ejecta, complementing the spatial analysis provided by the maps.
@@
According to Fig. 6, although the O density is significantly lower than in the prominent dense shell shown in Figs. 2 and 3, O-rich ejecta are also present in the innermost regions of the remnant, which are rich in Fe and other heavy elements (as evident from Fig. 4). Lowering the density threshold for selecting cells substantially increases the filling factor, reaching approximately 0.5 at a 1% threshold and about 0.8 at a 0.1% threshold of the maximum O-rich ejecta density (see upper panels in the figure). Although web-like structures and filaments rich in O and Fe can also be seen in the innermost ejecta (Figs. 4 and 6), these features may appear a bit less sharp and less concentrated than in the outer regions. This visual effect may be connected to differences in the relative spatial resolution $\delta r/r$ when a uniform cartesian grid with cell size $\delta x$ is applied to resolve structures at different radii $r$. In Appendix A, we evaluated the impact of spatial resolution on structures forming at different distances $r$ from the explosion center. It may be possible that filaments forming in the innermost regions of the remnant are resolved with lower accuracy, leading to slightly smoother and more diffuse structures, as observed in the simulations.
@@
The presence of O-rich ejecta in the remnant’s interior results in a LoS velocity distribution characterized by a prominent peak at $u_{\rm LoS}\approx-1300$ km s-1, consistent with expectations from Fig. 5. Moreover, the velocity distribution extends almost continuously up to $u_{\rm LoS}\approx 4500$ km s-1, exhibiting a broad range of velocities between $-2500$ and $4500$ km s-1. We also observe a pronounced asymmetry in the velocity distribution of unshocked ejecta, with material moving toward us reaching maximum absolute velocities of $\approx-2500$ km s-1, while ejecta traveling away exhibit maximum absolute velocities of about $\approx+4500$ km s-1. This result is consistent with the analysis of Spitzer observations of Cas A (Isensee et al. 2010). The extended asymmetric velocity distribution underscores the complex kinematics of the O-rich ejecta and highlights their dynamic mixing within the remnant’s interior.
@@ 3.2 Generation of the web-like network of ejecta filaments
Figure 11 shows the distributions of Ni- and O-rich ejecta roughly one day after the shock breakout, approximately 22 hours after the core collapse. The densest regions of Ni-rich ejecta, with densities exceeding $2\times 10^{-8}$ g cm-3, are displayed as a red isosurface in the left panels. At this stage, the web-like network of O-rich ejecta filaments is already well developed (see left panels), showcasing its intricate complexity. According to the evolution before the shock breakout described above, this structure reflects the turbulent mixing induced by the expansion of bubbles of neutrino-heated material and the effects of HD instabilities (see also Wongwathanarat et al. 2015). In other words, the network of filaments underscores the dynamic interplay of neutrino-driven convection, which accelerates the efficient mixing of heavy elements as the blast wave propagates outward through the progenitor star’s layers. Additionally, the figure reveals notable inversions of the ejecta layers in specific regions, consistent with observations of similar phenomena in Cas A, as discussed in previous studies (Orlando et al. 2016, 2021). For instance, in the extended Ni-rich plume expanding eastward, O-rich ejecta are found to occupy a physically interior position relative to Ni-rich material. Such layer inversions are indicative of the highly asymmetric and turbulent nature of the explosion, driven by the initial anisotropies in the neutrino heating and further amplified during the expansion.
@@ 4 Summary and Conclusions
We have investigated the origin and evolution of the intricate network of ejecta filaments recently unveiled in Cas A by JWST, using high-resolution 3D HD and MHD simulations. Since the inclusion of magnetic fields did not significantly alter our results (see Appendix B), we focused here on the HD simulations. By tracing the evolution of the remnant from the core-collapse SN explosion to an age of $\approx 1000$ years, our study sheds light on the physical processes that can shape a filamentary structure of ) ejecta in young SNRs.
Our simulations, based on a neutrino-driven SN explosion model, demonstrate that a web-like network of ejecta filaments forms naturally during the early phases of the remnant evolution, shaped by a combination of key processes. These include the expansion of neutrino-heated bubbles immediately following the core collapse, HD instabilities developing as the blast wave propagates through the stellar interior, and the Ni-bubble effect, which becomes significant shortly after the shock breakout. This latter effect enhances mixing and compresses lighter elements along the boundaries of initially Ni-rich domains, further contributing to the formation of a filamentary structure. The result is a dense, interconnected web of filaments enriched in O and other intermediate-mass elements (e.g., C, Ne, Mg, Si).
@@
  • Filamentary structure at the age of Cas A. The filamentary network, dominated by O-rich ejecta, forms an interconnected web-like structure that closely matches the morphology observed by JWST, although smaller length-scale features observed in Cas A are not present, most likely due to the effect of numerical diffusion on features occupying a few cells. The modeled filaments exhibit distinct spatial organization, with intermediate-mass and light elements (e.g., C, O, Ne, Mg) forming the network, while heavier elements (e.g., Ca, Ti, and Fe) are more centrally concentrated or confined to dense plumes. The filaments align systematically with channels on the surface of high-density Fe-rich ejecta, similar to rivers filling valleys. O-, Ne-, and Mg-rich ejecta permeate these channels, creating the intricate web-like pattern of ejecta observed by JWST. Large voids are present within the network, particularly in regions disrupted by extended Fe-rich plumes, consistent with ejecta layer inversions observed in the Fe-rich regions of Cas A. The model accurately reproduces the observed filament thickness down to scales of $\sim 0.01$ pc, emphasizing the role of early explosion dynamics.

### v2

@@ 1 Introduction
In this paper, we investigate this issue using the 3D HD/MHD model from Orlando et al. (2021), complemented by high-resolution simulations designed to capture ejecta features at scales approaching those observed by JWST. This model traces the evolution of a neutrino-driven SN explosion, replicating key characteristics of Cas A from the immediate post-core-collapse phase to its current state. In such a way, this approach provides a robust framework to link the observed filaments with specific physical processes active throughout the SN and the remnant evolution. Our study focuses on two main objectives: (i) characterizing the morphological and kinematic properties of the network of filament by comparing the simulated structures with the results from observations, (ii) tracing the ejecta evolution from the initial seconds after the explosion to examine how filamentary features observed today are connected to physical processes and phenomena triggered shortly after core collapse.
@@
The paper is organized as follows: Sect. 2 summarizes our HD/MHD simulation setup and methodology; Sect. 3 presents the results, focusing on the 3D structure of the unshocked ejecta and their origins tracing back to the core-collapse; finally, Sect. 4 summarizes our main conclusions and discusses their broader implications for understanding SN explosion mechanisms and remnant evolution. In Appendix A, we investigate the effects of the spatial resolution on the results; in Appendix B, we evaluate the effects of the magnetic field on the structure of unshocked ejecta; in Appendix C, we describe the content of online multi-media material, offering interactive and dynamic representations of the phenomena discussed in the paper.
2 The 3D HD/MHD model
@@
The output from the 3D SN simulation at $\approx 17.85$ hours after core-collapse served as the initial condition for 3D simulations that track the transition from the SN phase to the SNR phase and the subsequent expansion of the remnant through the CSM (Orlando et al. 2021, 2022). The CSM environment is modeled as a stellar wind with a density profile proportional to $r^{-2}$, normalized to $0.8$ cm-3 at a radius of 2.5 pc (in agreement with Lee et al. 2014), and includes an asymmetric, dense circumstellar shell with its densest region positioned on the near side to the northwest (Orlando et al. 2022). This shell is included to account for significant asymmetries of the reverse shock, which cannot be explained by models assuming a spherically symmetric stellar wind from the progenitor (see Orlando et al. 2022). Recent JWST observations support the existence of this shell, having revealed evidence of a shocked, dense CSM structure around Cas A (the so-called “Green Monster”), consistent with past interactions between the remnant and a dense region of CSM material (Milisavljevic et al. 2024; De Looze et al. 2024; Orlando et al. 2025). In the MHD simulations, the ambient magnetic field is represented by a “Parker spiral” configuration, resulting from the progenitor star’s rotation and the corresponding outward flow of the stellar wind (Parker 1958; see Orlando et al. 2019a for further details).
The SNR simulations include: (i) energy deposition from radioactive decay, specifically the 56Ni $\rightarrow$ 56Co $\rightarrow$ 56Fe chain, modeled by an internal energy term that assumes local deposition, excluding neutrinos that escape freely, and assuming no $\gamma$-ray leakage from the inner part of the remnant (Orlando et al. 2021); (ii) deviations from ionization equilibrium, estimated through the maximum ionization age in each cell of the spatial domain (Orlando et al. 2015); and (iii) deviations from electron-proton temperature equilibration, with initial electron heating at shock fronts up to $kT=0.3$ keV (Ghavamian et al. 2007), and post-shock ion/electron temperatures determined via Coulomb collisions (Orlando et al. 2015). In this paper, we omitted the impact of cosmic ray back-reaction at shock fronts explored in previous studies (Orlando et al. 2016, 2022). This process primarily influences the dynamics of the forward shock and the structure of the mixing region between the forward and reverse shocks (e.g., Orlando et al. 2012; Pavlović et al. 2018). However, it does not significantly affect the evolution of the unshocked ejecta, which is the primary focus of this work.
@@ 3 Results
Table 2: Masses of ejecta elements (total, shocked, and unshocked) in the W15-IIb-sh-HD+dec-hr model at 351 years.
@@
Notes: The table also includes the ratio of shocked to unshocked masses. The last three columns provide, for comparison, the shocked mass values inferred from Chandra data analysis of Cas A (Hwang & Laming 2012) and the unshocked mass values inferred from Spitzer data analysis (Laming & Temim 2020). $(a)$ Data from Hwang & Laming (2012); each element is presented with two values, corresponding to assumed filling factors based on density spike thickness of $2.5^{\prime\prime}$ and $5^{\prime\prime}$; $(b)$ Data from Laming & Temim (2020); $(c)$ The ejecta mass of 44Ti is overestimated by a factor $10-20$ due to a reduced network in the original SN model, which included only 9 of 13 $\alpha$ nuclei up to 56Ni (see Wongwathanarat et al. 2017); $(d)$ The values reported consider the total Fe masses from incomplete Si burning (FeSi) plus from complete Si burning and/or $\alpha$-rich freezout (Feα) (Hwang & Laming 2012); $(e)$ The tracer nucleus 56X represents Fe-group species synthesized in neutron-rich environments as, for instance, in neutrino-heated ejecta (see Wongwathanarat et al. 2017). The table lists the mass of 56Fe, calculated as $M_{\rm Fe}+0.5M_{\rm X}$, and the mass of 56X, estimated as $0.5M_{\rm X}$.
@@ 3.1 The structure of unshocked ejecta at the age of Cas A
From our analysis of the distribution of unshocked species, we found no significant differences between models with or without magnetic fields (see Table 1). Therefore, in the following, we focused on the results of the run W15-IIb-sh-HD+dec-hr, noting any differences observed in the magnetized case (W15-IIb-sh-MHD+dec-hr) where relevant. In Appendix B, we report a selection of results from run W15-IIb-sh-MHD+dec-hr. The effects of magnetic fields, particularly in shaping the structure of the Green Monster, are discussed in a separate study (Orlando et al. 2025). Unless otherwise stated, the time in this study is measured from the core bounce (see Wongwathanarat et al. 2017), also referred to as the time after core-collapse.
@@
Figure 6: Maps of the filling factor of O-rich ejecta, calculated for cells where the O density exceeds thresholds of 0.1% (upper panel), 1% (middle panel), or 10% (bottom panel) of the maximum O-rich ejecta density in the unshocked volume. The maps assume the model orientation matching Cas A.
Figure 7: Normalized LoS velocity distributions corresponding to the maps of the filling factor of O-rich ejecta shown in Fig. 6, derived from the same cells used to derive the filling factor maps.
@@
To investigate further this issue and to facilitate a more quantitative comparison between model predictions and JWST observations, we calculated the filling factor of O-rich ejecta and the corresponding LoS velocity distributions. Figure 6 displays maps of the filling factor, assuming the model is oriented to match Cas A. These maps represent the filling factors computed along the LoS for cells where the O density exceeds specific thresholds ($f_{\rm ths}=$ 0.1%, 1%, or 10%) of the maximum O-rich ejecta density within the entire unshocked volume. Specifically, for each LoS, the filling factor is calculated as the ratio of the volume of cells containing O-rich ejecta above a given density threshold ($V_{\rm O}$) to the total volume of all cells containing unshocked ejecta along the same LoS ($V_{\rm tot}$). This ratio provides a normalized measure of the extent to which the unshocked ejecta volume is occupied by O-rich material, offering valuable insights into its spatial distribution and density contrasts. Figure 7 shows the corresponding normalized LoS velocity distributions. These distributions are derived from the same cells used to compute the filling factors and highlight the kinematic properties of the O-rich ejecta, complementing the spatial analysis provided by the maps.
@@
According to Fig. 6, although the O density is significantly lower than in the prominent dense shell shown in Figs. 2 and 3, O-rich ejecta are also present in the innermost regions of the remnant, which are rich in Fe and other heavy elements (as evident from Fig. 4). Lowering the density threshold for selecting cells substantially increases the filling factor, reaching approximately 0.5 at a 1% threshold and about 0.8 at a 0.1% threshold of the maximum O-rich ejecta density (see upper and middle panels in the figure). Although web-like structures and filaments rich in O and Fe can also be seen in the innermost ejecta (Figs. 4 and 6), these features may appear a bit less sharp and less concentrated than in the outer regions. This visual effect may be connected to differences in the relative spatial resolution $\delta r/r$ when a uniform cartesian grid with cell size $\delta x$ is applied to resolve structures at different radii $r$. In Appendix A, we evaluated the impact of spatial resolution on structures forming at different distances $r$ from the explosion center. It may be possible that filaments forming in the innermost regions of the remnant are resolved with lower accuracy, leading to slightly smoother and more diffuse structures, as observed in the simulations.
@@
The presence of O-rich ejecta in the remnant’s interior results in a LoS velocity distribution characterized by a prominent peak at $u_{\rm LoS}\approx-1300$ km s-1 (see Fig. 7), consistent with expectations from Fig. 5. Moreover, the velocity distribution extends almost continuously up to $u_{\rm LoS}\approx 4500$ km s-1, exhibiting a broad range of velocities between $-2500$ and $4500$ km s-1. We also observe a pronounced asymmetry in the velocity distribution of unshocked ejecta, with material moving toward us reaching maximum absolute velocities of $\approx-2500$ km s-1, while ejecta traveling away exhibit maximum absolute velocities of about $\approx+4500$ km s-1. This result is consistent with the analysis of Spitzer observations of Cas A (Isensee et al. 2010). The extended asymmetric velocity distribution underscores the complex kinematics of the O-rich ejecta and highlights their dynamic mixing within the remnant’s interior.
@@ 3.2 Generation of the web-like network of ejecta filaments
Figure 12 shows the distributions of Ni- and O-rich ejecta roughly one day after the shock breakout, approximately 22 hours after the core collapse. The densest regions of Ni-rich ejecta, with densities exceeding $2\times 10^{-8}$ g cm-3, are displayed as a red isosurface in the left panels. At this stage, the web-like network of O-rich ejecta filaments is already well developed (see left panels), showcasing its intricate complexity. According to the evolution before the shock breakout described above, this structure reflects the turbulent mixing induced by the expansion of bubbles of neutrino-heated material and the effects of HD instabilities (see also Wongwathanarat et al. 2015; Vartanyan et al. 2024). In other words, the network of filaments underscores the dynamic interplay of neutrino-driven convection, which accelerates the efficient mixing of heavy elements as the blast wave propagates outward through the progenitor star’s layers. Additionally, the figure reveals notable inversions of the ejecta layers in specific regions, consistent with observations of similar phenomena in Cas A, as discussed in previous studies (Orlando et al. 2016, 2021). For instance, in the extended Ni-rich plume expanding eastward, O-rich ejecta are found to occupy a physically interior position relative to Ni-rich material. Such layer inversions are indicative of the highly asymmetric and turbulent nature of the explosion, driven by the initial anisotropies in the neutrino heating and further amplified during the expansion.
@@ 4 Summary and conclusions
We have investigated the origin and evolution of the intricate network of ejecta filaments recently unveiled in Cas A by JWST, using high-resolution 3D HD and MHD simulations. Since the inclusion of magnetic fields did not significantly alter our results (see Appendix B), we focused here on the HD simulations. By tracing the evolution of the remnant from the core-collapse SN explosion to an age of $\approx 1000$ years, our study sheds light on the physical processes that can shape a filamentary structure of unshocked ejecta in young SNRs.
Our simulations, based on a neutrino-driven SN explosion model, demonstrate that a web-like network of ejecta filaments forms naturally during the early phases of the remnant evolution, shaped by a combination of key processes. These include the expansion of neutrino-heated bubbles immediately following the core collapse, HD instabilities developing as the blast wave propagates through the stellar interior, and the Ni-bubble effect, which becomes significant shortly after the shock breakout. This latter effect enhances mixing and compresses lighter elements along the boundaries of initially Ni-rich domains, further contributing to the formation of a filamentary structure. The result is a dense, interconnected web of filaments enriched in O and other intermediate-mass elements (e.g., C, Ne, Mg, and Si).
@@
  • Filamentary structure at the age of Cas A. The filamentary network, dominated by O-rich ejecta, forms an interconnected web-like structure that closely matches the morphology observed by JWST, although smaller length-scale features observed in Cas A are not present, most likely due to the effect of numerical diffusion on features occupying a few cells. The modeled filaments exhibit distinct spatial organization, with intermediate-mass and light elements (e.g., C, O, Ne, and Mg) forming the network, while heavier elements (e.g., Ca, Ti, and Fe) are more centrally concentrated or confined to dense plumes. The filaments align systematically with channels on the surface of high-density Fe-rich ejecta, similar to rivers filling valleys. O-, Ne-, and Mg-rich ejecta permeate these channels, creating the intricate web-like pattern of ejecta observed by JWST. Large voids are present within the network, particularly in regions disrupted by extended Fe-rich plumes, consistent with ejecta layer inversions observed in the Fe-rich regions of Cas A. The model accurately reproduces the observed filament thickness down to scales of $\sim 0.01$ pc, emphasizing the role of early explosion dynamics.